Aspen Company estimates its manufacturing overhead to be $1,170,000 and its direct labor costs to be $600,000 for year 2. Aspen worked on three jobs for the year. Job 2-1, which was sold during year 2, had actual direct labor costs of $180,000. Job 2-2, which was completed, but not sold at the end of the year, had actual direct labor costs of $305,000. Job 2-3, which is still in work-in-process inventory, had actual direct labor costs of $130,000. Actual manufacturing overhead for year 2 was $1,200,000. Manufacturing overhead is applied on the basis of direct labor costs Required a. How much overhead was applied to each job in year 2? Job 2-1 Job 2-2 Job 2-3 b. What was the over- or underapplied manufacturing overhead for year 2? manufacturing overhead
Answer #1

a)

Statement of Overhead Applied:

Direct labor costs

Weight of Direct labor cost

Overhead Applied

Job 2-1

$180000

$180000/$615000 = 0.29

$1200000*0.29 = $348000

Job 2-2

$305000

$305000/$615000 = 0.50

$1200000*0.50 = $600000

Job 2-3

$130000

$130000/$615000 = 0.21

$1200000*0.21 = $252000

TOTAL

$615000

Actual manufacturing overhead

$1200000

b)

Actual manufacturing overhead = $1200000

Estimated manufacturing overhead = $1170000

Over applied overhead for Year 2 = Actual manufacturing overhead - Estimated manufacturing overhead

= $1200000 - $1170000 = $30000